A very strong, but inept, shot putter puts the shot straight up vertically with an initial speed of 10.6 m/s.


1
Expert's answer
2021-09-12T19:00:45-0400

Given:

"v_0=10.6\;\rm m/s"


The maximum height the shot reaches

"h_{\max}=\frac{v_0^2}{2g}=\frac{10.6^2}{2*9.8}=5.73\:\rm m"